OpenStudy (osanseviero):

I shoot something vertically. The speed was 25m/s when the object was at 1/4 of the max height. Which was the initial speed?

OpenStudy (shamim):

Let the maximum height is h

OpenStudy (shamim):

So u will get 2 equations 25^2=u^2-2*9.8*h/4 And 0^2=u^2-2*9.8*h

OpenStudy (shamim):

Frm above 2 equations find u=?

OpenStudy (shamim):

I used equation v^2=u^2-2*g*h